A firm is considering a new project that will generate cash revenue of $1,300,000 and cash expenses of $700,000 per year for five years. The equipment necessary for the project will cost $300,000 and will be depreciated straight-line over four years. What is the NPV if the firm's marginal tax rate is 35% and discount rate is 10%?
Answer #1

depreciation under SLM method = cost / number of years

= 300000 / 4 =$75,000

free cash flows = (revenue - expenses - depreciation)*(1 - tax) + depreciation

so free cash flows for the first four years = (1,300,000 - 700,000 - 75,000)*(1 - 0.35) + 75000

=$416,250

free cash flow for the fifth year = (1,300,000 - 700000 - 0) * (1 - 0.35)

= $390,000 (no depreciation in the last year)

present value of cash flows:

NPV = present value of cash flows - initial cash outflow

NPV = 1,561,615.8 - 300,000

= $1,261,615.8